# A Self-Improving Coding Agent

### v1

@@ Abstract
We demonstrate that an LLM coding agent, equipped with basic coding tools, can autonomously edit itself, and thereby improve its performance on benchmark tasks. We find performance gains from 17% to 53% on a random subset of SWE Bench Verified, with additional performance gains on LiveCodeBench, as well as synthetically generated agent benchmarks. Our work represents an advancement in the automated and open-ended design of agentic systems, and provides a reference agent framework for those seeking to post-train LLMs on tool use and other agentic tasks.
@@ 1 Introduction
LLMs have recently made impressive advancements across a range of domains and tasks (Anthropic, ; Google, 2024; OpenAI, 2024). However, in order to put these LLMs to use in real world applications, LLMs must be wrapped in code to orchestrate them and expose ‘tools’ that allow the models to take actions. These action-taking LLMs are referred to as agents, and the broader system an agent system.
These agent systems often show dramatic improvements in benchmark performance over “plain” LLMs (Yao et al., 2023b; Zelikman et al., 2023; Chen et al., 2023), through combinations of prompting strategies and methods for combining different LLM outputs. Early examples include best-of-$N$ sampling and simple prompting strategies such as chain of thought (Kojima et al., 2022). However more sophisticated schemes have shown success in getting the desired behavior and performance improvements from the models, for instance STaR (Zelikman et al., 2022), Tree of Thoughts (Yao et al., 2023a), Graph of Thoughtsagent (Besta et al., 2023), LLM Debate (Du et al., 2023), Iterative Self-Refinement (Madaan et al., 2023), Expert Prompting (Long et al., 2024) among many others. The comprehensive survey of Schulhoff et al. (2024) demonstrates the vast number of manually created strategies to date.
Recent improvements in coding agents raise the question of whether these agents themselves can autonomously modify and improve their own code by discovering e.g. new prompting schemes or tools without manual design and implementation. We argue that this style of fully self-referential meta-agent programming is possible today and offers a sound alternative to the ad-hoc, trial-and-error approach of hand-crafted orchestrators, which may only explore a small fraction of the solution space. Recent work in the Automated Design of Agentic Systems (ADAS) (Hu et al., 2024) uses a meta-agent to optimise agent implementations. However, Hu et al. (2024) is not self improving, as there are two separate agents: the target-agent that performs the task, and the meta-agent, which improves the target agent.
@@
  • Empirical evidence that self-referential agents can effectively improve their own implementations; we find performance improves from 17% to 53% performance on a random subset of SWE Bench Verified, even with consideration given to safety constraints and resource efficiency.
@@
We make our code available at https://github.com/MaximeRobeyns/self_improving_coding_agent.
@@ 2 Related Work
The traditional approach to developing and optimizing agent systems has been to manually design agent architectures and prompting techniques. Notable examples include Chain-of-Thought prompting (Wei et al., 2022), self-refinement (Madaan et al., 2023) and self-reflection (Shinn et al., 2023) for improving reasoning, tool use frameworks (Schick et al., 2023), and various compositional agent systems (Wang et al., 2023; Ahn et al., 2022). While these hand-crafted approaches have achieved strong results, they require significant human effort and may miss useful patterns that could be discovered through automated search.
@@
Another direction focuses on enabling agents to learn reusable skills and continuously self-improve. MaestroMotif (Klissarov et al., 2024) uses LLM feedback to learn skill rewards and combines skills through code generation. This builds on earlier work on intrinsically motivated reinforcement learning (Chentanez et al., 2004) and autotelic agents (Colas et al., 2022) that develop repertoires of internally motivated skills, as well as work in open-endedness Zhang et al. (2024); Faldor et al. (2024) that use LLMs to identify interesting and useful directions to explore in.
@@
Several approaches leverage LLMs to optimize agent behaviors through natural language interaction. OPRO (Yang et al., 2024) and Promptbreeder (Fernando et al., 2023) focus on optimizing prompts through language. Others have explored using LLMs to critique and refine agent behaviors (Klissarov et al., 2023), generate training curricula (Kumar et al., 2024), or provide natural language feedback for reinforcement learning (Qu et al., 2024). Since our agent can edit its entire codebase, this includes the ability to tune its own prompts to optimise the behaviour of any part of the agent.
@@
Recent work has begun exploring automated approaches for designing and optimizing agent systems. AgentSquare (Shang et al., 2024) proposed a modular design space that abstracts agent components into planning, reasoning, tool use and memory modules, allowing automated search over module combinations.
@@
Perhaps the most closely related line of prior work began with ADAS (Hu et al., 2024). ADAS used a target-agent which performs the actual task, and a meta-agent which improves the target-agent. As such, ADAS is not self-improving (as the meta-agent improves the target agent, not itself). Moreover, in ADAS the meta-agent edits only a single forward function, written in a domain-specific language which has been carefully designed to make expressing different prompting schemes very straightforward. In contrast, our self-improving coding agent is fully self-improving (i.e. there is no distinction between the meta and target agent), and it operates over the agent’s full codebase.
@@
Of course, we would expect the first truly self-improving agents to be coding agents, because agents are written in code. The natural approach is to start off with a basic coding agent that can open/close/edit files, run commands in the terminal etc, then to launch this agent in a self-improvement loop. We believe that our self-improving coding agent is the first such work. However, there are two papers claiming self-improving agents, but they do not evaluate in the coding setting, as they do not consider “full” coding agents. First, Gödel Agent (Yin et al., 2024) has specific tools (such as action_adjust_logic and action_read_logic) that allow modification of small parts of the agent as it is running. Thus, it is not a general-purpose coding agent, as it is traditionally understood. And as such, as with ADAS, it was evaluated on language understanding and mathematical benchmarks (DROP Dua et al., 2019, MGSM Shi et al., 2022, MMLU Hendrycks et al., 2020 and GPQA Rein et al., 2023), rather than coding benchmarks. Second, Zelikman et al. (2024) introduce a self-taught optimizer for recursively self-improving code generation. Again, this is not a general-purpose coding agent with the ability to e.g. work on arbitrary software engineering tasks, and as such, it self-improves and evaluates and only on algorithmic tasks such as learning parity with noise (Blum et al., 2003), String Grid Distance, and 3-SAT.
@@ 3 Methods
The main running loop of SICA resembles the ADAS loop Hu et al. (2024). In particular, both SICA, and ADAS keep an archive of previous agents and their benchmark results. In the first step, SICA takes the best performing agent from the archive up until this point as the meta-agent, which is instructed to look through the archive, identify an improvement, and implement it. The high-level scheme is illustrated in Algorithm 1. Note that this is different from ADAS, in that ADAS has a fixed meta-agent, so does not select the meta-agent from the archive (the archive in ADAS is an archive of target-agents, not meta-agents). Second, both ADAS and SICA evaluate the new agent on a set of evaluation benchmarks and store the results.
@@
We begin by describing the initial coding agent. We then go on to describe the benchmark runner framework, and how this framework naturally allows us to create a self-referential task (that of improving the coding agent.
@@ 3.1 Initial agent
We were interested in useful, practically relevant coding agents were capable of self-improvement. To that end, we started with a sophisticated coding agent, with a number of tools, sub-agents and oversight mechanisms. Of course, the tools, sub-agents and oversight mechanisms were all amenable to self-improvement; but here we simply describe the starting point.
@@
First, we designed a set of tools to open (i.e. add to the context) and close files, overwrite files, execute a shell command and a simple calculator tool. We also include basic tools to submit a final answer to a benchmark problem, return a result from a sub-agent call to its caller, as well as exiting early from a sub-agent call. Finally, we provide an ‘archive analysis’ tool, which facilitates extracting high-level summary statistics about each agent iteration in the archive. We describe the tool calling mechanism in more detail in Appendix C.
@@
Next, we included several sub-agents, which are invoked much like a tool call by accepting input arguments and returning a string value. Introducing this abstraction serves two purposes. The first is to ensure the benchmark tasks can run entirely autonomously, without human direction and guidance. We provide the problem statement to the main entrypoint agent, (the full prompt for which is given in Appendix A) which then decomposes the task, delegating and asking questions to sub-agents. The second is to manage the context length for extended agent runs. The meta-improvement task is challenging, and the main agent may take on the order of one hour to finish one meta-improvement iteration. The sub-agent calls allow the context used for a sub-task to be naturally discarded when the agent returns, with the value returned by the sub-agent providing a summary of that sub-task in the main agent’s context. Initially, we included three agents: a coding agent, a problem-solver agent and a reasoning agent. The full prompts for these agents are given in Appendix A.1.
@@
Finally, one important component that enables extended, autonomous operation while remaining on-task, avoiding loops or other behaviours that frustrate progress is the asynchronous overseer. This is an LLM, running in a concurrent thread, which is prompted to judge whether there are any pathological behaviours or deviations from the original task, and which may intervene by notifying the currently running agent (or any of its parents) by means of a message in-lined into the assistant context, or in serious cases to cancel the execution of an agent if it has deviated too far or appears stuck.
@@
This overseer is provided with a comprehensive textual representation of the state of the system (a truncated example of which is provided in Appendix B). This consists of the ‘callgraph’ of agent and sub-agent execution, in addition to the event stream of LLM messages, tool calls and responses, and sub-agent calls and responses associated with each agent call. This broad representation of the system’s behaviour, allows the overseer to to pick out patterns, as well as detect repeated work completed by previous agent calls. The full Overseer prompt is given in Appendix A.2.
@@ 4 Experiments and Results
We first show a run that includes relatively long-horizon tasks in the benchmark set. We include SWE Bench Verified (Chowdhury et al., 2024) which tests problem decomposition, code navigation as well as fast and efficient file editing from which we select a fixed random subset of 50 questions. We also include 50 random questions from LiveCodeBench (Jain et al., 2024), the questions from which resemble competitive programming and often include more theoretical reasoning.
@@
We run the agent system using Sonnet 3.5 (v2) (Anthropic, ) in most of the agents, with the exception of a ‘reasoning agent’ which uses o3 mini (OpenAI, 2025). The results are shown in Figure 3, with more detailed iteration-level peformance numbers in Table 1.
Table 1: Performance per iteration: the benchmark accuracy numbers correspond to a synthetic file editing benchmark and symbol location (codebase navigation) benchmark, SWE Bench Verified, and LiveCodeBench, respectively. The metrics are averaged across all problem instances across all benchmarks.
@@ 4.1 Performance on Reasoning Tasks
We also evaluate the effectiveness of the self-referential agent system in improving task performance in more reasoning-heavy domains. We ran another experiment with two question answering tasks in the benchmark set: AIME 2024 and GPQA Diamond Rein et al. (2023), the results of which are shown in Figure 4.
@@
Inspecting the traces, we observe that for many of the runs, the main agent merely delegated the problem to the o3-mini based reasoning agent, and did not leverage the full capabilites of the agent system. For ‘reasoning models’ such as o1 (OpenAI, 2024), o3-mini or DeepSeek’s R1 (DeepSeek-AI, 2025)), we suspect that the inclusion of crude components aiming to induce reasoning behaviour (such as those included in iterations 4 and 6 in Figure 4) may in fact interrupt the reasoning chain of thought of a reasoning model trained outside the agent system, resulting in a drop in performance. We look forward to future work training ‘agent’ LLMs jointly with the design of the agent system which will hopefully make better use of both components’ strengths.
@@ 5 Conclusion
We provide SICA, a self-improving coding agent. Our results demonstrate that self-improvement, implemented via an agent system editing its own codebase, is indeed possible, particularly in longer-horizon ‘agentic’ tasks where the base LLMs stand to gain more from the structure and guidance of a scaffolding system—for instance showing improvements from 17% to 53% on a random subset of SWE Bench Verified, while also achieving a slight decrease in the average time spent per problem on average. The framework we present offers both a practical approach to automated agent system design and a foundation for future work in jointly training foundation models with their agent systems.

### v2

@@ Abstract
Recent advancements in Large Language Models (LLMs) have spurred interest in deploying LLM agents to undertake tasks in the world. LLMs are often deployed in agent systems: code that orchestrates LLM calls and provides them with tools. We demonstrate that an agent system, equipped with basic coding tools, can autonomously edit itself, and thereby improve its performance on benchmark tasks. We find performance gains from 17% to 53% on a random subset of SWE Bench Verified, with additional performance gains on LiveCodeBench, as well as synthetically generated agent benchmarks. Our work represents an advancement in the automated and open-ended design of agentic systems, and demonstrates a data-efficient, non gradient-based learning mechanism driven by LLM reflection and code updates.
@@ 1 Introduction
LLMs have recently made impressive advancements across a range of domains and tasks [2, 14, 29]. However, in order to put these LLMs to use in real world applications, LLMs must be wrapped in code to orchestrate them and expose tools that allow the models to take actions. These action-taking LLMs are referred to as agents, and the broader system an agent system.
These agent systems often show dramatic improvements in benchmark performance over “plain” LLMs [43, 46, 5], through combinations of prompting strategies and methods for combining different LLM outputs. Early examples include best-of-$N$ sampling and simple prompting strategies such as chain of thought [20]. However more sophisticated schemes have shown success in getting the desired behavior and performance improvements from the models, for instance STaR [45], Tree of Thoughts [42], Graph of Thoughts [3], LLM Debate [10], Iterative Self-Refinement [23], Expert Prompting [22] among many others. The comprehensive survey of Schulhoff et al. [34] demonstrates the vast number of manually created strategies to date.
Recent improvements in coding agents raise the question of whether these agents themselves can autonomously modify and improve their own code by discovering e.g. new prompting schemes or tools without manual design and implementation. We argue that this style of fully self-referential meta-agent programming is possible today and offers a sound alternative to the ad-hoc, trial-and-error approach of hand-crafted orchestrators, which may only explore a small fraction of the solution space. Recent work in the Automated Design of Agentic Systems (ADAS) [16] uses a meta-agent to optimise agent implementations. However, Hu et al. [16] is not self-improving, as there are two separate agents: the target-agent that performs the task, and the meta-agent, which improves the target agent. A motivation for a self-improving system is that the improvements in coding abilities may be leveraged during subsequent improvement steps, hopefully compounding.
@@
  • Empirical evidence that self-referential agents can effectively improve their own implementations; we find performance improves from 17% to 53% performance on a random subset of SWE-Bench Verified, even with consideration given to safety constraints and resource efficiency.
@@ 2 Related Work
The traditional approach to developing and optimizing agent systems has been to manually design agent architectures and prompting techniques. Notable examples include Chain-of-Thought prompting [39], self-refinement [23] and self-reflection [37] for improving reasoning, tool use frameworks [33], and various compositional agent systems [38, 1]. While these hand-crafted approaches have achieved strong results, they require significant human effort and may miss useful patterns that could be discovered through automated search.
@@
Another direction focuses on enabling agents to learn reusable skills and continuously self-improve. MaestroMotif [19] uses LLM feedback to learn skill rewards and combines skills through code generation. This builds on earlier work on intrinsically motivated reinforcement learning [6] and autotelic agents [8] that develop repertoires of internally motivated skills, as well as work in open-endedness [48, 12] that use LLMs to identify interesting and useful directions to explore in.
@@
Several approaches leverage LLMs to optimize agent behaviors through natural language interaction. OPRO [41] and Promptbreeder [13] focus on optimizing prompts through language. Others have explored using LLMs to critique and refine agent behaviors [18], generate training curricula [21], or provide natural language feedback for reinforcement learning [31]. Since our agent can edit its entire codebase, this includes the ability to tune its own prompts to optimise the behaviour of any part of the agent.
@@
Recent work has begun exploring automated approaches for designing and optimizing agent systems. AgentSquare [35] proposed a modular design space that abstracts agent components into planning, reasoning, tool use and memory modules, allowing automated search over module combinations. AlphaEvolve [28] set out to use a coding agent to make scientific discoveries and optimize computational infrastructure, the results of which may accelerate the training of the LLM underpinning AlphaEvolve itself. The approach taken by these works leans more heavily on structured, evolutionary search than our own.
@@
Perhaps the most closely related line of prior work began with ADAS [16]. ADAS used a target-agent which performs the actual task, and a meta-agent which improves the target-agent. As such, ADAS is not self-improving (as the meta-agent improves the target agent, not itself). Moreover, in ADAS the meta-agent edits only a single forward function, written in a domain-specific language which has been carefully designed to make expressing different prompting schemes very straightforward. In contrast, our self-improving coding agent is fully self-improving (i.e. there is no distinction between the meta and target agent), and it operates over the agent’s full Python codebase.
@@
Of course, we would expect the first truly self-improving agents to be coding agents, because agents are written in code. The natural approach is to start off with a basic coding agent that can open/close/edit files, run commands in the terminal etc, then to launch this agent in a self-improvement loop. We believe that our self-improving coding agent is the first such work. However, there are two papers claiming self-improving agents, but they do not evaluate in the coding setting, as they do not consider “full” coding agents. First, Gödel Agent [44] has specific tools (such as action_adjust_logic and action_read_logic) that allow modification of small parts of the agent as it is running. Thus, it is not a general-purpose coding agent, as it is traditionally understood. And as such, as with ADAS, it was evaluated on language understanding and mathematical benchmarks (DROP [11], MGSM [36], MMLU [15] and GPQA [32]), rather than coding benchmarks. Second, Zelikman et al. [47] introduce a self-taught optimizer for recursively self-improving code generation. Again, this is not a general-purpose coding agent with the ability to e.g. work on arbitrary software engineering tasks, and as such, it self-improves and evaluates and only on algorithmic tasks such as learning parity with noise [4], String Grid Distance, and 3-SAT.
@@ 3 Methods
The main running loop of SICA resembles the ADAS loop Hu et al. [16] and is illustrated in Figure 1. In particular, both SICA, and ADAS keep an archive of previous agents and their benchmark results. In the first step, SICA takes the best performing agent from the archive up until this point as the meta-agent, which is instructed to look through the archive, identify an improvement, and implement it. The high-level scheme is illustrated in Algorithm 1. Note that this is different from ADAS, in that ADAS has a fixed meta-agent, so does not select the meta-agent from the archive (the archive in ADAS is an archive of target-agents, not meta-agents). Second, both ADAS and SICA evaluate the new agent on a set of evaluation benchmarks and store the results.
@@
We begin by describing the initial coding agent. We then go on to describe the benchmark runner framework, and how this framework naturally allows us to create a self-referential task (that of improving the coding agent).
@@ 3.1 Initial agent
We were interested in useful, practically relevant coding agents capable of self-improvement. To that end, we started with a sophisticated coding agent, with a number of tools, sub-agents and oversight mechanisms. Of course, the tools, sub-agents and oversight mechanisms were all amenable to self-improvement; but here we simply describe the starting point.
@@
We designed a set of tools to open (i.e. add to the context) and close files, overwrite files, execute a shell command and a simple calculator tool. We also include basic tools to submit a final answer to a benchmark problem, return a result from a sub-agent call to its caller, as well as exiting early from a sub-agent call. Finally, we provide an ‘archive analysis’ tool, which facilitates extracting high-level summary statistics about each agent iteration in the archive. We describe the tool calling mechanism in more detail in Appendix C.
@@
Next, we included several sub-agents. These are merely agents invoked by an existing agent, much like a tool call, by accepting input arguments and returning a string value. Introducing this abstraction serves two purposes. The first is to ensure the benchmark tasks can run entirely autonomously, without human direction and guidance. We provide the problem statement to the main entrypoint agent, (the full prompt for which is given in Appendix A) which then decomposes the task, delegating and asking questions to sub-agents. The second is to manage the context length for extended agent runs. The meta-improvement task is challenging, and the main agent may take on the order of one hour to finish one meta-improvement iteration. The sub-agent calls allow the context used for a sub-task to be naturally discarded when the agent returns, with the value returned by the sub-agent providing a summary of that sub-task in the main agent’s context. Initially, we included three agents: a coding agent, a problem-solver agent and a reasoning agent. The full prompts for these agents are given in Appendix A.1.
@@
Finally, one important component we provide in the initial agent that enables extended, autonomous operation while remaining on-task, avoiding loops or other behaviours that frustrate progress is the asynchronous overseer. This is an LLM, running periodically in a concurrent thread, which is prompted to judge whether there are any pathological behaviours or deviations from the original task, and which may intervene by notifying the currently running agent (or any of its parents) by means of a message in-lined into the assistant context, or in serious cases to cancel the execution of an agent if it has deviated too far or appears stuck. The overseer is called every 30s by default, or on key events.
@@
This overseer is provided with a comprehensive textual representation of the state of the system (a truncated example of which is provided in Appendix B) with verbose content truncated to limit context. This consists of the ‘callgraph’ of agent and sub-agent execution, in addition to the event stream of LLM messages, tool calls and responses, and sub-agent calls and responses associated with each agent call. This broad representation of the system’s behaviour, allows the overseer to to pick out patterns, as well as detect repeated work completed by previous agent calls. The full Overseer prompt is given in Appendix A.2.
@@ 4 Experiments and Results
We first show a run that includes relatively long-horizon tasks in the benchmark set. We include SWE Bench Verified [7] which tests problem decomposition, code navigation as well as fast and efficient file editing from which we select a fixed random subset of 50 questions. We also include 50 random questions from LiveCodeBench [26], the questions from which resemble competitive programming and often include more theoretical reasoning [^2] Both SWEBench-Verified and LiveCodeBench are MIT licensed.
@@
We run the agent system using Sonnet 3.5 (v2) [2] in most of the agents in the agent system, with the exception of a ‘reasoning agent’ which uses o3 mini [30]. While the base agent system can run with much smaller and less expensive models, and many of our benchmark tasks can be solved by this class of models, the meta-improvement task is a complicated long-horizon task which requires powerful models. The API cost of the 15 iteration run was approximately \$7,000.
The results are shown in Figure 3, with more detailed iteration-level peformance numbers in Table 1.
Table 1: Performance per iteration: the benchmark accuracy numbers correspond to a synthetic file editing benchmark and symbol location (codebase navigation) benchmark, SWE-Bench Verified, and LiveCodeBench, respectively. We also show the metrics averaged across all problem instances across all benchmarks to compare the relative ‘per problem’ costs across agent iterations.
@@ 4.1 Performance on Reasoning Tasks
We also evaluate the effectiveness of the self-referential agent system in improving task performance in more reasoning-heavy domains. We ran another experiment with two question answering tasks in the benchmark set: AIME 2024 and GPQA Diamond [32], the results of which are shown in Figure 4.
@@
Inspecting the traces, we observe that for many of the runs, the main agent merely delegated the problem to the o3-mini based reasoning agent, and did not leverage the full capabilites of the agent system. For ‘reasoning models’ such as o1 [29], o3-mini or DeepSeek’s R1 [9]), we suspect that the inclusion of crude components aiming to induce reasoning behaviour (such as those included in iterations 4 and 6 in Figure 4) may in fact interrupt the reasoning chain of thought of a reasoning model trained outside the agent system, resulting in a drop in performance. We look forward to future work training ‘agent’ LLMs jointly with the design of the agent system which will hopefully make better use of both components’ strengths.
@@ 5 Conclusion
We provide SICA, a self-improving coding agent. Our results demonstrate that self-improvement, implemented via an agent system editing its own codebase, is indeed possible, particularly in longer-horizon ‘agentic’ tasks where the base LLMs stand to gain more from the structure and guidance of a scaffolding system—for instance showing improvements from 17% to 53% on a random subset of SWE-Bench Verified, while also achieving a slight decrease in the average time spent per problem on average. The framework we present offers both a practical approach to automated agent system design and a foundation for future work in jointly training foundation models with their agent systems.
5.1 Limitations
Our initial attempt at a self-improving coding agent is not without limitations. One key difficulty was getting the LLM-based agent to autonomously come up with truely novel, innovative, feasible and interesting modification ideas at each meta-improvement step, which is a theme which has been commented on in the open-ended learning literature [27, 40, 17]. The cost of settling on a bad idea which suffered from poor ‘taste’ was a lengthy agent editing step followed by an even more expensive run through the benchmarks. While the failed iteration persists in the archive, in principle acting as an example of what not to do, we found that the initial feature ideas would often heavily influence later feature ideas as variations on the same theme. This path dependency may lead to higher variance agent runs; with poor quality initial feature suggestions (e.g. fixating on caching open files) often lowering the quality of subsequent feature suggestions.
We also note that in optimizing for agent running time and cost, our relatively short 5-minute timeouts (and to a lesser degree per-problem cost limits) cause the initial benchmark scores to perhaps be lower than expected for the underlying language model (e.g. Sonnet 3.5 v2), especially for longer-horizon benchmark tasks like SWE-Bench. Indeed, much of the initial improvement we report arises from features speeding up file editing and agent cost, which are not traditionally reflected to the same degree in SWE-Bench numbers.
5.2 Future Work
Our approach reflects a non weight-based learning mechanism, where we update the system’s behaviour through a combination of optimising the prompts on which the model is conditioned, as well as the way the LLM calls are orchestrated by editing the underlying code. This is in some sense very sample-efficient, where a detailed inspection of the archive of outcomes informs substantial modifications at each step. However, we recognize the potential of jointly updating the foundation model’s weights with the agent system. The work of AlphaEvolve [28] demonstrates a proof of this principle, and it may be fruitful to more explicitly fine-tune the foundation model to work with new newly created tools and scaffold at each iteration.
While we did not get to the point of benchmark saturation in our relatively short runs, we expect that future systems may run out of things to learn from a static benchmark set. One way this problem may be solved is through the automated design and curation of benchmarks, as initially demonstrated in the work of [12] where RL environments are designed in code. By including two synthetic benchmarks in our work which were curated entirely by the agent through manual prompting (the file editing and symbol navigation benchmark) we demonstrate that SICA is flexible enough to support creating new benchmarks and evaluations for itself.